ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: Xu, Zhong Ling, National Natural Science Foundation of China The Popularization of Mathematical Methods in China In the sixties, professor Hua Lao-Keng of the Chinese Academy of Sciences switched the main emphasis of his research work from pure to applied mathematics. Hua has developed a number of practical approaches, adapted to the peculiar conditions of china, for popularizing mathematical methods in china. The popularization of mathematical methods in china is commonly known as the popularization of “ optimization and overall planning ” or simply as the “two methods”. Optimization problems can be divided into three major categories; Problems without obvious mathematical expression, whose solutions have to be sought by “ trial ” methods. Problems with clear mathematical expressions, such as the linear and nonlinear programming problems. Problems related to time or dynamic problems. Professor Hua has found that the critical path method (CPM) is one of the most useful tools which can also be easily popularized in china. This application work began with the organization of young and energetic popularization teams. In 20 years of popularizing and applying the “ two methods ”, Prof. Hua and his teams visited twenty-two provinces, hundreds of cities, and thousands of factories, meeting millions of people.